Maricopa's Success scholarship fund receives a gift of $ 170000. The money is invested in stocks, bonds, and CDs. CDs pay 3.25 % interest, bonds pay 4.4 % interest, and stocks pay 11.4 % interest. Maricopa Success invests $ 25000 more in bonds than in CDs. If the annual income from the investments is $ 10055 , how much was invested in each account?
Maricopa Success invested $
in stocks.
Maricopa Success invested $
in bonds.
Maricopa Success invested $

Answer by stanbon(75887)

Maricopa's Success scholarship fund receives a gift of $ 170000. The money is invested in stocks, bonds, and CDs. CDs pay 3.25 % interest, bonds pay 4.4 % interest, and stocks pay 11.4 % interest. Maricopa Success invests $ 25000 more in bonds than in CDs. If the annual income from the investments is $ 10055 , how much was invested in each account?
--------------------------
Equations:
s + b + C = 170000
11.4s + 4.4b + 3.25C = 1005500
b = C + 25000
-----------------------
Rearrange::
s + b + C = 170000
1140s + 440b + 325C = 100550000
0 + b - C = 25000
--------------------------
Use any method you know to solve the system to get:
stocks = 45000
bonds = 75000
CD's = 50000
----------------
Cheers,
Stan H.
--------------


Maricopa Success invested $
in stocks.
Maricopa Success invested $
in bonds.
Maricopa Success invested $